OpenStudy (anonymous):

Evaluate each expression when x = 5 and y = -4. I'm going to draw the problem out.

OpenStudy (anonymous):



OpenStudy (samwisegamge):

https://mathway.com/

zepdrix (zepdrix):

\[\large\rm \frac{\sqrt x}{y}+\frac{y}{x}\quad=\quad \frac{\sqrt5}{-4}+\frac{-4}{5}\]What do you mean by evaluate? Would they like you to combine the fractions?

OpenStudy (anonymous):

I'm not sure what they mean but I just know the answer to it. I just don't know how to get to the answer.

OpenStudy (jjcoolja12):

well i guess the answer is −1.35901699…

OpenStudy (anonymous):